Session: Oct 2008 (session 3)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (A. Cattani at Tue Oct 28 17:22:11 2008) Minor failures :
We have lost data from 16:45 UT to the end. Switeched to a bad disk bank

Shanghai (Bo Xia at Tue Oct 28 20:34:15 2008) Minor failures : TSYS
There was something wrong with the cryogenic LNA of L-band,so our engineer had repaired the cryogenic LAN. After UT 8:34,the receiver was okay. VSN:NAIC-003/2000/1024

Urumqi (Wenjun Yang at Wed Oct 29 00:41:50 2008) Success :

Torun (Kaz Borkowski at Wed Oct 29 07:17:53 2008) Minor failures :
At 17:52 UT the telescope took its round trip in the azimuth (unwrapping cables), so the record is off source for about 10 minutes. No other known problems.

Effelsberg (Uwe Bach at Wed Oct 29 09:07:28 2008) Minor failures : FS
Lost some hours due to problems with the antenna control computer: 2:08 to 4:32 UT, some minutes ~7 UT, 12:35 to 12:48 UT, and some minutes around 13:50 UT.

Onsala (Michael Lindvist at Wed Oct 29 09:56:15 2008) Minor failures : WIND
RFI. Late start due to high winds (rain as well...). Started around 03:40 UTC. Antenna wrap around 19:30 UTC. Otherwise no known problems.

Westerbork (Geert Kuper at Wed Oct 29 13:34:10 2008) Success :
Wb used in phased array mode. No known problems

Robledo (Cristina Garcia-Miro at Thu Oct 30 15:39:22 2008) Minor failures : RFI
L-band RFI impacted VC03 USB during whole observation. Station time delay: +0.7109 microsec.

Noto (P. Cassaro at Mon Nov 3 13:32:58 2008) Success :